# Mitigating Multimodal Hallucinations via Gradient-based Self-Reflection

### v1

@@ Abstract
Hallucinations in multimodal large language model are caused by the text-visual bias and the co-occurrence bias. The former reflects an over-reliance on text information in the decision-making process, while the latter arises from the statistical object-pairing patterns abstracted from the training data. Existing mitigation methods heuristically address these biases without understanding the fluctuating bias level across the instances. We first propose estimating the influence of respective token types (visual, prompt, and previous outputs) using a gradient-based self-reflection method. The estimated token influence further enables the detection of object-related visual tokens and their integration into an influence-aware contrastive decoding framework to mitigate both types of biases simultaneously. Our method operates without the need for additional resources, such as costly fine-tuning, extra models, or data statistics. Extensive experiments show it effectively reduces hallucinations, achieving up to a $92\%$ accuracy increase on LLaVA-QA90.
@@ 1 Introduction
Figure 1: Overview of our method, which acts as a token-wise post-processing mechanism to mitigate text-visual hallucinations by enhancing visual influence (colored bars) in prediction. Our token-influence approach explicitly aligns this amplification with the most influential text inputs—either prompts (gray bar) or previous outputs (white bar). By leveraging token-wise influence measurement, the method also mitigates co-occurrence bias by selectively avoiding the enhancement of visual tokens (green and magenta bars) linked to previously mentioned objects.
@@
Multimodal Large Language Models (MLLMs) liu2024llava ; dai2023instructblipgeneralpurposevisionlanguagemodels ; ye2024mplug ; bai2023qwen ; chen2024internvl , have recently gained significant momentum in generating coherent, context-aware content across diverse domains. Despite their impressive capabilities, these models face a significant challenge: hallucination, where the generated text content is not grounded in the visual inputs rohrbach2018object ; li2023evaluating . This issue undermines the trust in the outputs of MLLMs.
Hallucinations in MLLMs are found to arise from two key biases studied in the literature li2023evaluating ; kang2023impact ; kim2024discovering : text-visual bias, and co-occurrence bias. Text-visual bias arises when a model relies too heavily on textual input, and neglects the visual information. This bias is also linked to the generated text length zhou2023analyzing ; favero2024multi . Longer text increases the risk of hallucinations, often stemming from early predictions that overly influence subsequent content generation. Co-occurrence bias arises from statistical correlations in the training data li2023evaluating , causing models to incorrectly predict objects based on their frequent pairings even when those objects are absent from the visual input. This bias is challenging to address, and currently lacks mitigation solutions that are free from statistical awareness.
Inference-based hallucination mitigation methods leng2024mitigating ; chen2024halc receive more attention than training-based approaches chen2023mitigating ; jiang2024hallucination ; yue2024less ; ben2023mocha ; sun2023aligning ; kang2023impact due to their cost effectiveness, as they eliminate the necessity for data collection, statistical analysis, and model training. Existing inference-related methods leng2024mitigating ; park2024convis ; favero2024multi ; woo2024don mitigate hallucination heuristically but lack a clear understanding of bias severity. Among inference-related methods, revision approaches xing2024efuf ; deng2024seeing ; radford2021learning often require additional models for prediction validation, such as MLLM-based scoring, which may also hallucinate. Meanwhile, image-level methods leng2024mitigating ; favero2024multi ; zhao2024mitigating fail to address co-occurrence hallucinations at the token (individual object) level. Existing token-level approaches woo2024don ; kan2024catchcomplementaryadaptivetokenlevel ; chen2024halc either lack semantic awareness of visual tokens or rely on auxiliary segmentation or detection models to obtain decoupled visual representations. Therefore, a method that jointly and quantitatively addresses all biases without requiring additional data, or models remains a challenging and unresolved issue.
In this work, to quantitatively estimate biases, we introduce a gradient-based token influence, which leverages the first-order Taylor expansion spivak1980calculus to estimate how each input token influences the output. This token influence enables the detection of the corresponding visual token of each predicted object (noun). In addition, we incorporate both token influence and the detected object-related visual tokens into a contrastive decoding framework li2022contrastive and propose Gradient-based Influence-Aware Contrastive Decoding (GACD) to jointly mitigate text-visual and co-occurrence biases. To address text-visual bias, GACD enhances visual token influence, aligning it with the most influential text components from prompts or previous outputs. This ensures visual grounding answer and prevents visual forgetting. To mitigate the co-occurrence bias, GACD selectively avoids enhancing visual tokens linked to objects mentioned in previous outputs. As illustrated in Fig. 1, GACD corrects the hallucinated prediction of ‘beer’ by enhancing the influence of visual tokens unrelated to ‘fork’ or ‘spoon’, resulting in a more visually grounded output. As text-visual bias tends to occur in long sentence generation zhou2023analyzing , we further introduce a sample-dependent stopping criterion based on visual influence to explicitly capture visual contributions and account for length effects.
Contributions of this work is summarized as:
  • We propose GACD, a novel hallucination mitigation method that reduces text-visual bias by enhancing the influence of visual tokens, without requiring external models or fine-tuning. Its token-wise adjustment further enables the mitigation of co-occurrence bias without relying on training dataset statistics.
  • We propose token influence to express MLLM biases, providing a detailed understanding at both the token (individual object) level and the sample level.
@@
  • Extensive experiments demonstrate that GACD effectively mitigates hallucinations and enhances accuracy without sacrificing information. GACD achieves up to $8\%$ increase in overall score on AMBER wang2023llm , an $8\%$ F1 boost on POPE li2023evaluating , up to $45\%$ improvement in detailness and a $92\%$ accuracy gain on LLaVA-QA90 liu2024visual .
@@ 2 Related Work
Hallucination and Bias. Hallucinations in LLMs often arise from biases in the training data mckenna2023sources ; huang2025survey , while in MLLMs, studies tonmoy2024comprehensive ; li2023evaluating ; fu2024mmecomprehensiveevaluationbenchmark show that hallucinations are closely linked to biases like text-visual and co-occurrence biases. Additionally, biases related to output position, which increase the risk of hallucination as output length grows, have been examined in favero2024multi ; zhou2023analyzing . Existing methods li2023evaluating ; fu2024mmecomprehensiveevaluationbenchmark ; kim2024discovering typically report only overall statistics, lacking a mathematical, sample-wise bias measurement. This distinction is important, as biases can vary case by case. Our approach measures sample-dependent bias via token-level gradient sensitivities, revealing how pre-trained MLLM parameters embed these biases  kim2019learning ; guo2024bias , and enabling self-reflective hallucination mitigation.
Hallucination Mitigation. Training-related hallucination mitigation methods chen2023minigpt ; jiang2024hallucination ; yue2024less ; ben2023mocha ; sun2023aligning ; lee-etal-2022-deduplicating are expensive, requiring access to training data and specialized statistical analysis to effectively mitigate hallucinations. Reinforcement-learning approaches martino2023knowledge ; xing2024efuf ; deng2024seeing ; radford2021learning ; zhai2024hallecontrolcontrollingobjecthallucination rely on supplementary feedback, often from human annotators or auxiliary LLMs/MLLMs, and the latter may themselves hallucinate. By contrast, post-decoding techniques modify model logits at inference time without further training or external feedback, making them lightweight add-ons. In text-only LLMs, such methods aim to align outputs with factual knowledge chuang2023dola ; lee2022factuality ; li2023inference . In MLLMs, post-decoding strategies emphasize the role of visual inputs li2023inference ; leng2024mitigating ; zhao2024mitigating ; favero2024multi and can be classified into image-level and token-level interventions. Image-level decoding methods leng2024mitigating ; zhao2024mitigating ; favero2024multi ; zhang2024debiasing ; manevich2024mitigatinghallucinationslargevisionlanguage treat all objects in the input image uniformly, limiting their effectiveness in addressing co-occurrence hallucinations. Existing token-level methods either rely on external segmentation chen2024halc and detection models kan2024catchcomplementaryadaptivetokenlevel or lack awareness of object-related decoupling woo2024don . In contrast, our method estimates embedded bias and decouples object-aware visual tokens, enabling comprehensive and sample-specific hallucination mitigation without relying on external data or models.
@@ 3 Method
In this section, we provide background on MLLMs, introduce the concept of token influence, and explain how GACD balances token influence through contrastive decoding to mitigate hallucinations.
@@ 3.1 Background on MLLMs
Consider a sequence $\mathbf{t}^{p}=[t^{p}_{1},\dots,t^{p}_{N}]$ as the input prompt, where $t^{p}_{n}$ ($1\leq n\leq N$) is a prompt token from a predefined vocabulary. Visual tokens $\mathbf{t}^{v}=[t^{v}_{1},\dots,t^{v}_{S}]$ are extracted from visual inputs $V$. The MLLMs generate the response sequence $\mathbf{y}=[y_{1},\dots,y_{M}]$ using the logit generation function $\mathcal{F}_{\theta}(\cdot)$ and the softmax function $\sigma(\cdot)$ 111 Probability of softmax output is interpreted as confidence. as:
$\mathbf{y}=\sigma(\mathcal{F}_{\theta}(\mathbf{t}^{v},\mathbf{t}^{p})).$ (1)
$y_{m}$ denotes an individual output token for $1\leq m\leq M$. The conditional probability distribution $p(\mathbf{y}|\mathcal{F}_{\theta},\mathbf{t}^{p},\mathbf{t}^{v})$ can therefore be expressed as:
$p(\mathbf{y}|\mathcal{F}_{\theta},\mathbf{t}^{v},\mathbf{t}^{p})=\prod_{m=1}^{M}p(y_{m}|\mathcal{F}_{\theta},\mathbf{t}^{v},\mathbf{t}^{p},\mathbf{y}_{<m}),$ (2)
where $\mathbf{y}_{<m}=[y_{1},\dots,y_{m-1}]$ for $m>1$ and is empty for $m=1$. Pre-trained MLLMs encode not only factual knowledge but also statistical correlations and biases from the training data kim2019learning ; guo2024bias . These biases are encoded within the model parameters ($\theta^{\star}$), enabling self-reflective bias interpretation.
@@ 3.2 Gradient-Based Token Influence Estimation
To capture these learned biases, we measure the impact each individual token has on the output logits. To achieve this, we propose token influence estimation based on a first-order Taylor expansion spivak1980calculus on logit generation function $\mathcal{F}_{\theta^{*}}$  [^2] Taylor expansion details are in supplementary Sec. A.:
$\mathcal{F}_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m}\approx\sum^{S}_{s=1}\mathbf{g}^{v}_{ms}{t}^{v}_{s}+\sum^{N}_{n=1}\mathbf{g}^{p}_{mn}{t}^{p}_{n}+\sum^{m-1}_{i=1}\mathbf{g}^{y}_{mi}y_{i}+Const,$ (3)
where
$\mathbf{g}^{v}_{ms}:=\frac{\partial(\mathcal{F}_{\theta^{*}})_{m}}{\partial{t}^{v}_{s}}\bigg{|}_{\mathbf{t}^{v}=\mathbf{t}^{v(0)}},\quad\mathbf{g}^{p}_{mn}:=\frac{\partial(\mathcal{F}_{\theta^{*}})_{m}}{\partial{t}^{p}_{n}}\bigg{|}_{\mathbf{t}^{p}=\mathbf{t}^{p(0)}},\quad\mathbf{g}^{y}_{mi}:=\frac{\partial(\mathcal{F}_{\theta^{*}})_{m}}{\partial y_{i}},$ (4)
are the first order gradients w.r.t., visual tokens $\mathbf{t}^{v}$, text prompt $\mathbf{t}^{p}$ and previous outputs $y_{i}$ and $Const$ denotes all other terms that are constant w.r.t., the $\mathbf{t}^{v},\mathbf{t}^{p}$.
We measure the importance of input tokens using the norm of their gradients. Specifically, we use the Manhattan (L1) norm, as it effectively captures the contribution of individual tokens. Accordingly, we represent the influence of each token as $\big{|}\mathbf{g}^{v/p/y}_{m\cdot}\big{|}$. The overall influence of visual tokens, text prompts, and previous outputs can be represented as:
$\texttt{I}^{v}_{m}=\sum^{S}_{s=1}\big{|}\mathbf{g}^{v}_{ms}\big{|},\quad\texttt{I}^{p}_{m}=\sum^{N}_{n=1}\big{|}\mathbf{g}^{p}_{mn}\big{|},\quad\texttt{I}^{y}_{m}=\sum^{m-1}_{i=1}\big{|}\mathbf{g}^{y}_{mi}\big{|}.$ (5)
This expression allows us to interpret bias at the individual sample level. We also define the influence of each visual token on a specific output vocabulary token $c$ as:
$I_{ms}(c)=\big{|}\mathbf{g}^{v}_{ms}[c]\big{|},$ (6)
@@
where $\mathbf{g}^{v}_{ms}[c]$ represents the gradient from the output vocabulary $c$ with respect to the input tokens.
3.3 Influence-Aware Contrastive Decoding
We extend contrastive decoding li2022contrastive by incorporating object-aware negative guidance and dynamically computing decoding weights based on token influence. As shown in Fig. 2, for each $m^{th}$ output, contrastive logits are generated alongside the original logits, with the decoding weight $\alpha_{m}$ inferred based on token influence from both.
Figure 2: Overview of the proposed GACD method. GACD adjusts contrastive decoding based on token influence (Sec. 3.2), explicitly amplifying the influence of visual tokens $\mathbf{t}^{u}$. The contrastive logits, acting as negative guidance, are generated by object-related visual token $\mathbf{t}^{o}$, prompt token $\mathbf{t}^{p}$, and previous output token $\mathbf{t}^{y}$, with object-related visual token $\mathbf{t}^{o}$ included only for noun predictions to reduce the influence of potential anchor objects and mitigate co-occurrence bias. The decoding weight $\alpha_{m}$ is calculated based on token influence in both original and contrastive logits, amplifying object-unrelated visual token $\hat{\mathtt{I}}^{u}_{m}$ to align with the dominant text component $\max({\color[rgb]{0,0,1}\hat{\mathtt{I}}^{p}_{m}},{\color[rgb]{0.0,0.5,0.0}\hat{\mathtt{I}}^{y}_{m}})$. For noun predictions, GACD detects the highest-influence visual token using individual visual influence $I_{ms}(y_{m})$ and flags it as object-related $\mathbf{t}^{o}$ for subsequent processing.
We begin with object-related visual token detection. To identify objects in previous outputs $\mathbf{y}_{<m}$, we first perform noun detection, treating each noun as an object. For each identified object, a mask is applied to flag the highest-influence object-related visual tokens. The mask $\mathcal{M}_{ms}$ fuses visual tokens related to all previously mentioned objects:

### v2

@@ Abstract
Multimodal large language models achieve strong performance across diverse tasks but remain prone to hallucinations, where outputs are not grounded in visual inputs. This issue can be attributed to two main biases: text–visual bias, the overreliance on prompts and prior outputs, and co-occurrence bias, spurious correlations between frequently paired objects. We propose Gradient-based Influence-Aware Constrained Decoding (GACD), an inference-based method, that addresses both biases without auxiliary models, and is readily applicable to existing models without finetuning. The core of our approach is bias estimation, which uses first-order Taylor gradients to understand the contribution of individual tokens—visual features and text tokens—to the current output. Based on this analysis, GACD mitigates hallucinations through two components: (1) suppressing spurious visual features correlated with the output objects, and (2) rebalancing cross-modal contributions by strengthening visual features relative to text. Experiments across multiple benchmarks demonstrate that GACD effectively reduces hallucinations and improves the visual grounding of MLLM outputs.
@@ 1 Introduction
Figure 1: Overview of our influence-aware constrained decoding framework, which mitigates hallucinations by regulating token-level influence. It reduces text–visual bias by enhancing visual token influence (blue bars) in alignment with the most influential text inputs— prompts (gray) or previous outputs (white). It further mitigates co-occurrence bias through anchor-specific suppression, selectively suppressing visual tokens (green, magenta) anchored to previously emitted nouns.
@@
Recent advances in Multimodal Large Language Models (MLLMs) have demonstrated remarkable capabilities in producing coherent and context-aware content across a wide range of domains (Bai et al., 2023; Dai et al., 2023; Chen et al., 2024b; Liu et al., 2024a; Ye et al., 2024). Despite their impressive advancements, these models remain prone to hallucination, wherein the generated text is not faithfully grounded in the visual modality (Rohrbach et al., 2018; Li et al., 2023b). This limitation poses a critical barrier to establishing trust in the outputs of MLLMs.
The hallucinations observed in MLLMs can be largely attributed to two fundamental biases (Kang & Choi, 2023; Li et al., 2023b; Kim et al., 2024). Text-visual bias refers to the excessive reliance on textual information—such as the input prompt and previously generated outputs—while neglecting the visual modality during generation. This bias becomes particularly pronounced in longer sequences, where MLLMs tend to depend more heavily on prior text and increasingly disregard visual cues (Zhou et al., 2023b; Favero et al., 2024). Co-occurrence bias arises from spurious statistical correlations embedded in the training data, which lead models to erroneously predict the presence of non-existent objects based on their frequent co-occurrence with observed objects in the visual inputs (Li et al., 2023b). This bias is particularly challenging to mitigate, and existing approaches largely rely on statistical priors rather than offering statistically agnostic solutions (Kang & Choi, 2023; Zhou et al., 2023a).
Existing efforts to mitigate hallucinations in MLLMs can be broadly categorized into inference-based methods, which operate at the decoding stage (Chen et al., 2024a; Favero et al., 2024; Leng et al., 2024; Park et al., 2024; Woo et al., 2024), and training-based approaches, which intervene during model optimization (Ben-Kish et al., 2023; Chen et al., 2023b; Kang & Choi, 2023; Sun et al., 2023; Jiang et al., 2024; Yue et al., 2024). Inference-based approaches are valued for their cost-effectiveness, as they avoid the need for additional data collection, data bias examination, or extensive model retraining. However, these methods offer limited insight into the severity of underlying biases, leaving the root causes of hallucination insufficiently understood. In addition, some inference-based methods rely on auxiliary models—such as segmentation networks  Chen et al. (2024a), detection systems  Kan et al. (2024) , or even additional MLLMs (Radford et al., 2021; Deng et al., 2024; Xing et al., 2024)—which undermine their purported cost-effectiveness.
Another limitation of existing methods lies in their lack of granularity when adjusting the underlying biases in MLLMs. Most approaches rely on heuristically tuned priors, which vary across datasets and fail to generalize reliably Leng et al. (2024); Zhao et al. (2024). Moreover, they apply uniform weighting across all visual features, offering no mechanism to selectively adjust bias at the level of individual featuresZhang et al. (2024b); Manevich & Tsarfaty (2024). This coarse treatment limits their effectiveness in mitigating co-occurrence bias, which arises from spurious statistical correlations between objects that are often represented by distinct visual features.
In this work, we propose an inference-based method that simultaneously addresses both text–visual bias and co-occurrence bias, without relying on auxiliary models or external supervision. The core of our approach is the estimation of underlying bias, achieved by quantifying the contribution of individual tokens—both visual features and text tokens—through gradients derived from a first-order Taylor expansion. Building on this analysis, the method mitigates hallucinations by reweighting tokens via two key components: (1) suppressing the influence of visual features that exhibit strong spurious correlations with the current output token, thereby reducing co-occurrence bias; and (2) rebalancing cross-modal contributions by enhancing the role of visual features to align more closely with that of text tokens in generating the current output. As illustrated in Fig. 1, our method, GACD, corrects hallucinated predictions—such as the spurious generation of “beer” in the presence of “fork” and “spoon”—by amplifying the contributions of visual tokens unrelated to those nouns, leading to outputs that are more faithfully grounded in the visual modality. Note also that our method is readily applicable to existing MLLMs at inference time.
We summarize our main contributions as follows.
  • We introduce an inference-based method for hallucination mitigation in MLLMs, built on a principled estimation of underlying bias via gradients obtained from a first-order Taylor expansion. This estimation provides a mechanism for understanding and granularly adjusting their influences of individual visual features and text tokens on the generation of the current output token, all without requiring auxiliary models or finetuning.
  • We design two complementary modules: (i) suppression of spurious visual features correlated with the current output token to alleviate co-occurrence bias, and (ii) cross-modal rebalancing to enhance the contributions of visual features relative to text tokens, thereby addressing text–visual bias.
@@
  • Extensive experiments demonstrate that GACD mitigates hallucinations and enhances accuracy without sacrificing information. GACD achieves up to $8\%$ increase in overall score on AMBER Wang et al. (2023), an $8\%$ F1 boost on POPE Li et al. (2023b), up to $45\%$ improvement in detailness and a $92\%$ accuracy gain on LLaVA-QA90 Liu et al. (2024b).
@@ 2 Related Work
Hallucination and Bias. Hallucinations in LLMs often arise from biases in the training data McKenna et al. (2023); Huang et al. (2025), while in MLLMs, studies Tonmoy et al. (2024); Li et al. (2023b); Fu et al. (2024) show that hallucinations are closely linked to biases like text-visual and co-occurrence biases. Additionally, biases related to output position, which increase the risk of hallucination as output length grows, have been examined in Favero et al. (2024); Zhou et al. (2023b). Existing methods Li et al. (2023b); Fu et al. (2024); Kim et al. (2024) typically report only overall statistics, lacking a mathematical, sample-wise bias measurement. This distinction is important, as biases can vary case by case. Our approach measures sample-dependent bias via token-level gradient sensitivities, revealing how pre-trained MLLM parameters embed these biases  Kim et al. (2019); Guo et al. (2024), and enabling self-reflective hallucination mitigation.
Hallucination Mitigation. Training-related hallucination mitigation methods Chen et al. (2023a); Jiang et al. (2024); Yue et al. (2024); Peng et al. (2025); Zadeh et al. (2025) are expensive, requiring access to training data and specialized statistical analysis. Among them, LPOI Zadeh et al. (2025) also employs an object-aware framework, highlighting the effectiveness of modeling object-level information for mitigating hallucinations. Reinforcement‐learning approaches Xing et al. (2024); Deng et al. (2024); Zhai et al. (2024) rely on supplementary feedback, often from human annotators or auxiliary LLMs/MLLMs, and the latter may themselves hallucinate. By contrast, post-decoding techniques modify model logits at inference time without further training or external feedback, making them lightweight add-ons. In text‐only LLMs, such methods aim to align outputs with factual knowledge Chuang et al. (2023); Li et al. (2023a). In MLLMs, post-decoding strategies emphasize the role of visual inputs Leng et al. (2024); Zhao et al. (2024); Favero et al. (2024) and can be classified into image-level and token-level interventions. Image-level decoding methods Zhang et al. (2024b); Manevich & Tsarfaty (2024) treat all objects in the input image uniformly, limiting their effectiveness in addressing co-occurrence hallucinations. Existing token-level methods either rely on external segmentation Chen et al. (2024a) and detection models Kan et al. (2024) or lack awareness of object-related decoupling Woo et al. (2024). Moreover, these methods typically introduce an implicit trade-off between accuracy and informativeness, reducing hallucinations at the expense of omitting valid details. Attention-based methods Tang et al. (2025); Zhang et al. (2024a) require careful selection of specific layers and often introduce model-specific adjustments or heuristics. In contrast, our GACD directly estimates embedded bias and decouples object-aware visual tokens, enabling sample-specific hallucination mitigation without external data, models, or model-specific adjustments, while achieving a more favorable balance between accuracy and informativeness.
@@ 3 Method
In this section, we provide background on MLLMs, introduce the concept of token influence, and explain how GACD balances token influence to mitigate hallucinations.
@@ 3.1 Background on MLLMs
MLLMs generate a finite token sequence $\mathbf{y}=[y_{1},\dots,y_{M}]$ in response to a visual input (image or video) and a textual prompt. Let $\mathcal{V}$ be a finite vocabulary. The prompt is tokenized as $\mathbf{t}^{p}=[t^{p}_{1},\dots,t^{p}_{N}]$ with $t^{p}_{n}\in\mathcal{V}$. The visual input is encoded by a visual encoder into features, which are then projected into the token-embedding space $\mathbb{R}^{d}$, yielding visual tokens $\mathbf{t}^{v}=[t^{v}_{1},\dots,t^{v}_{S}]$ with $t^{v}_{s}\in\mathbb{R}^{d}$, where $d$ is the shared token embedding dimension used for $\mathcal{V}$.
A MLLM with parameters $\theta$ computes, at each decoding step $m$, a logit vector
$\mathbf{z}_{m}\;=\;\pi_{\theta}\!\left(\mathbf{t}^{v},\,\mathbf{t}^{p},\,\mathbf{y}_{<m}\right)\in\mathbb{R}^{|\mathcal{V}|},\qquad\mathbf{y}_{<m}=[y_{1},\dots,y_{m-1}]\text{ (empty when }m=1\text{)}.$ (1)
This induces a categorical next-token distribution via the softmax $\sigma:\mathbb{R}^{|\mathcal{V}|}\!\to\!\Delta^{|\mathcal{V}|-1}$:
$p_{\theta}\!\left(y_{m}\mid\mathbf{t}^{v},\mathbf{t}^{p},\mathbf{y}_{<m}\right)\;=\;\big[\sigma(\mathbf{z}_{m})\big]_{y_{m}},\qquad 1\leq m\leq M,$ (2)
where $\sigma(\mathbf{z}_{m})\in\Delta^{|\mathcal{V}|-1}$ denotes the probability distribution [^1] We use “confidence” to denote the model-assigned probability of the emitted token. over the vocabulary, and $\cdot]_{y_{m}}$ selects the component corresponding to token $y_{m}\in\mathcal{V}$. At inference, $y_{m}$ is sampled from this categorical distribution (e.g., greedy, beam search). The sequence likelihood factorizes by the chain rule:
$p_{\theta}\!\left(\mathbf{y}\mid\mathbf{t}^{v},\mathbf{t}^{p}\right)\;=\;\prod_{m=1}^{M}p_{\theta}\!\left(y_{m}\mid\mathbf{t}^{v},\mathbf{t}^{p},\mathbf{y}_{<m}\right).$ (3)
Given a dataset $\mathcal{D}$ of $(\mathbf{t}^{v},\mathbf{t}^{p},\mathbf{y})$, maximum-likelihood training (or fine-tuning) estimates $\theta^{\star}$ by maximizing the conditional log-likelihood. Pretrained MLLMs encode statistical regularities (including spurious correlations) from training data in $\theta^{\star}$; such behavior can be probed without changing $\theta^{\star}$ via parameter-dependent analyses (e.g., gradients/attributions or counterfactual decodings) Kim et al. (2019); Guo et al. (2024), enabling self-reflective bias interpretation.
@@ 3.2 Gradient-Based Token Influence Estimation
To capture these embeded biases, we measure how each input token perturbs the output logits. Let $\mathbf{z}^{\star}_{m}\in\mathbb{R}^{|\mathcal{V}|}$ denote the step-$m$ logits $\mathbf{z}^{\star}_{m}=\pi_{\theta^{\star}}(\mathbf{t}^{v},\mathbf{t}^{p},\mathbf{y}_{<m})$. Around a reference sample point $(\mathbf{t}^{v(0)},\mathbf{t}^{p(0)},\mathbf{y}_{<m}^{(0)})$, the first-order Taylor expansion Spivak (1980) of the logits $\mathbf{z}^{\star}_{m}$ is
$\mathbf{z}^{\star}_{m}\;\approx\;\sum^{S}_{s=1}\mathbf{g}^{v}_{ms}{t}^{v}_{s}+\sum^{N}_{n=1}\mathbf{g}^{p}_{mn}{t}^{p}_{n}+\sum^{m-1}_{i=1}\mathbf{g}^{y}_{mi}y_{i}+Const,$ (4)
where $Const$ denotes other terms that are constant w.r.t., $\mathbf{t}^{v}$ and $\mathbf{t}^{p}$ and the token-wise Jacobians are
$\mathbf{g}^{v}_{ms}\;\coloneqq\;\left.\frac{\partial\,\mathbf{z}^{\star}_{m}}{\partial\,\mathbf{t}^{v}_{s}}\right|_{\mathbf{t}^{v}=\mathbf{t}^{v(0)}},\qquad\mathbf{g}^{p}_{mn}\;\coloneqq\;\left.\frac{\partial\,\mathbf{z}^{\star}_{m}}{\partial\,\mathbf{t}^{p}_{n}}\right|_{\mathbf{t}^{p}=\mathbf{t}^{p(0)}},\qquad\mathbf{g}^{y}_{mi}\;\coloneqq\;\frac{\partial\,\mathbf{z}^{\star}_{m}}{\partial\,\mathbf{y}_{i}}\Big|_{\mathbf{y}=\mathbf{y}_{<m}^{(0)}},$ (5)
where $\big|_{\cdot}$ indicate evaluation at the reference sample point. Taylor expansion details are in supplementary Sec. A. Each $\mathbf{g}^{v}_{ms},\,\mathbf{g}^{p}_{mn},\,\mathbf{g}^{y}_{mi}$ indicate a small token perturbation in its corresponding embedding space to a perturbation of the predict logit vector in $\mathbb{R}^{|\mathcal{V}|}$. We score the importance of each input token by the Manhattan norm of its gradient:
$I^{v}_{ms}\;=\;\|\mathbf{g}^{v}_{ms}\|_{1},\qquad I^{p}_{mn}\;=\;\|\mathbf{g}^{p}_{mn}\|_{1},\qquad I^{y}_{mi}\;=\;\|\mathbf{g}^{y}_{mi}\|_{1},$ (6)
@@
and $I^{v}_{ms}[c]$ represents the gradient from the output vocabulary $c$ with respect to each visual tokens. Aggregating over tokens yields step-$m$ group-level influences:
$\texttt{I}^{v}_{m}\;=\;\sum_{s=1}^{S}I^{v}_{ms},\qquad\texttt{I}^{p}_{m}\;=\;\sum_{n=1}^{N}I^{p}_{mn},\qquad\texttt{I}^{y}_{m}\;=\;\sum_{i=1}^{m-1}I^{y}_{mi}.$ (7)
These quantities decompose, at the sample level, how visual tokens, prompt tokens, and prior outputs contribute to the logit of $y_{m}$, enabling interpretation of bias per sample.
3.3 Influence-Aware Constrained Decoding
GACD builds on token influence estimation with two components: (i) Object-aware Visual Token Grouping and (ii) Anchor-specific Influence-weighted Decoding. At step $m$, the former partitions visual tokens into object-related $\mathbf{t}^{o}$ and unrelated $\mathbf{t}^{u}$ based on objects detected in $\mathbf{y}_{<m}$. The latter extends contrastive decoding Li et al. (2022) by forming Anchor-specific negative guidance logits from pre-mentioned objects and computing a decoding weight $\alpha_{m}$ from token-influence measurements.